A patient, who weighs 90 kg, is to be given dobutamine by iv infusion at a rate of 7.5 micrograms/kg/minute. If one 25 ml vial of dobutamine, which contains 17.5 mg/ml, is diluted to 2 litres with glucose 5% w/v solution, at what rate in ml/minute, should the infusion be set?
Answer #1

patient weight: 90kg

Drug rate: 7.5MCG/kg/min

That is

7.5×90/minutes

625MCG/minute

1mg=100MCG

So,

625mcg=6.25mg/minutes

1 vial= 25ml

1ml=17.5mg

1vial=17.5×25= 437.5mg

2vial=2×437.5= 875mg.

As per prescription:

Patient is getting: 6.25mg/ minutes

For one hour: 6.25mg×60= 375/hour

For 2 hours: 6.25mg×120=750/hour

For 2 hour 20minutes: 6.25mg×140= 875mg/140minutes

In 2litre or 2000ml= 875mg

1000ml= 437.5mg

1ml= 0.4375mg

15ml= 6.25mg

Patient is getting= 6.25mg/minutes

Answer: 6.25mg/minutes

Or 15ml / minute